31 December 2007

RHIC Startup:  Good running over the weekend.  The running sum of luminosity is now corrected for accidentals. 

Saturday morning had fills with very bad Blue lifetime, due to a noise source discovered ten days ago when looking at the LF Shottky signal.  It appears as if something is turning on and off, but what is it?  Thomas, Wolfram and Peter looked but were unable to find a source.  There is no intensity dependence.  IPMs show the same emittance growth.  Studies of the slow logs and pscompare showed no guilty party. 

At the end of last week, we started doing automatic orbit corrections every half hour during physics stores.  This improves the beam lifetime and reduces the experimental background each time. 

Christoph changed the chromaticity to increase transmission efficiency. 

PHENIX reports the North side backgrounds are getting worse.  They request a new collimation or tuning effort.  This is new because they have traditionally been troubled by the Blue beam hitting the South Muon ID counters.  PHENIX will want to reverse their magnet field on Wednesday since they will have accumulated half of their total luminosity for the run.  They also have a Pad Chamber problem with the electronics for PC1.  This will be addressed on the Wednesday Maintenance Day. 

STAR saw a low pressure front that set their TPC into Purge mode since it couldn't keep up with the dropping barometric pressure.  They are still struggling with the high level triggers for the FMS.  Other than that they have been accumulating good luminosity.  STAR plans to open their magnet on Wednesday.  It has a small water leak.  Fixing this will be an 8 hour job on the Wednesday Maintenance day. 

Access Controls will remove the EBIS chipmunk from the LINAC UPS, decoupling AGS running EBIS. 

Polarimeter work was done on the Blue electronics, cycling the power to the Blue and Yellow HV supplies and the DVM.  This fixed the communication problems we were having last week.  Yellow preamps/shapers will be installed during Wednesday's Maintenance Day.  Currently there is a calibration run underway to understand the three different electrostatic shielding designs being tested for the preamp housing.   AGS cold snake is being cooled down already.  LINAC sections are in preparation for the polarized proton running.  LINAC will be ready in ~two weeks.  AGS polarization work will need to start ~two weeks ahead of pp running.

APEX work begins at 8:00PM tomorrow night, continuing to 8:00AM? Wednesday morning when Maintenance Day begins.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020